Suppose that a game is to be played with a single die 

assumed fair. In this game a player wins $20 if a 2 turns up, 

$ 40 if a 4 turns up; loses $30 if a 6 turns up; while the 

player neither wins nor loses if any other faces turns up. 

Find the expected sum of money to be won.


1
Expert's answer
2022-04-12T05:49:08-0400

"E[X]=20\cdot (1/6)^2+40\cdot (1/6)^4-30\cdot (1/6)^6=\$\ 0.59"